# Introducing Spotlight: A Novel Approach for Generating Captivating Key Information from Documents

### v1

@@ 7 Conclusion
In this paper, we introduced the concept of “Spotlight,” a novel, concise representation of source documents that offers a diiferent perspective to conventional summaries and highlights. Our work identifies a set of unique characteristics that differentiate spotlights, emphasizing their ability to present key aspects of a document in a faithful yet compelling way, effectively capturing the reader’s attention as well as drawing their attention to the source document. To validate this concept, we prepared four carefully annotated datasets, which served as the basis for a comprehensive analysis. Through rigorous quantitative evaluations and human studies, we established nine key features that differentiate spotlights primarily in terms of coverage and presentation style. Building on these findings, we proposed a Direct Preference Optimization (DPO)-based fine-tuning algorithm to generate spotlights. Extensive evaluation using both traditional metrics and LLM-based assessments demonstrate the effectiveness of our approach. We believe that, similar to summarization, there is a significant potential for developing various types of spotlights — such as aspect-based, query-focused, and personalized versions — which will be the focus of our immediate future work.
@@
Our datasets are neither multilingual nor multimodal. So, we need to explore how state-of-the-art approaches can be utilized in multlingual and multimodal scenarios for spotlight generation - which we aim to do as a part of future work. We have also explored some edge cases and limitations of the model generated spotlights as in Appendix J.
@@ Appendix A Dataset
Figure 2: Examples of (Document, Spotlight) pairs from different datasets, An example from CSPubSum is given in Figure 1, The documents and the example spotlight of Research Presentation have been truncated in this visualization owing to their length
@@
Figure 3: News Headline and the Short Description taken as spotlight for the News dataset
@@ B.1 Readability
Table 10: Feature level Analysis of Spotlight (Spot) vs. Human Annotated Summary (Summ) for 60 articles of each datasets, For Entropy, Compactness, Flesch Kincaid, Gunning score, Dale-Chall and McLaughlin’s score lower values signifies better result, whereas for Skew50, Flesch Reading score, FORCAST and Interest-Generation higher values signifies a better result as discussed in Section 3, More Exploratory Results of Table 3
@@ B.2 Informativeness and Degree of Extraction
We analyze the number of sentences in the document that achieve the R1, R2, and RL score above a specified threshold th with at least one sentence in the summary or spotlight. This approach assumes that if a sentence meets the threshold, its information has been incorporated into the summary or spotlight - i.e. informativeness. However, sentences incorporate information in an abstractive fashion so the threshold th needs to be set at a relatively low value. At a value of $th=0.3$ our findings of different Rouge (1, 2 and L) scores at the Column ‘Informativeness’ in Table 11 and 13 respectively for human-generated and GPT-4 generated summaries, indicates that a larger number of sentences from the source document are incorporated into summaries compared to spotlights. It signifies that the summary is more informative than the spotlight.
However, when the threshold value is increased (at a value of $th=0.8$) — indicating a higher level of extractiveness — the trend reverses, with fewer sentences meeting the threshold in summaries compared to spotlights as shown in ‘Degree of Extraction’ column using different Rouge (1, 2 and L) scores in Table 11 and 13 respectively for human-generated and GPT-4 generated summaries. The increased extractiveness of the spotlight can also be attributed to its enhanced readability. This trend is consistent across different ROUGE metrics (R1, R2, RL), as shown in the mentioned Tables.
@@
B.3 Spotlight-Highlight: Distinct Features
To explore the distinctness between Spotlight and Highlights, we have performed a similar mini-story study to Section 3.2. Details in Appendix Section G.3. In our analysis of distinct features (as described in Section 3.2), we observed that spotlights consistently outperform highlights in conveying a mini-story narrative. Annotator evaluations indicate that spotlights were rated as having a more engaging narrative in 72% of cases for CSPubSum, 84% for Research Presentation, 80% for News, and 72% for Wikipedia. While both spotlights and highlights aim to distill key information from source documents, spotlights uniquely present a self-contained narrative that not only summarizes but also sparks curiosity and encourages further exploration. This narrative element is less pronounced in traditional highlights, which tend to list isolated key points rather than weaving them into a cohesive mini-story.
B.4 Interest Generation aspects of a spotlight
@@
Figure 4: Reasoning on Interest-Generation aspect of a Spotlight on Life and Career of Anja Nissen from Wikipedia dataset
@@
Dataset Both of them Spotlight only Summary only None of them
News 6 34 6 4
CSPubSum 11 32 4 3
Wikipedia 12 31 1 6
Research 11 30 1 8
Presentation
@@ Appendix C Spotlight Generation Process
KTO Ethayarajh et al. (2024): We initially use the summaries for documents in the dataset to create a binary preference Kahneman-Tversky Optimization (KTO) dataset. Each document from the original dataset contributes one upvoted and one downvoted example. The document paired with the ground truth spotlight forms the upvoted example (indicating a preferred completion), while the document paired with a summary forms the downvoted example (denoting undesirable). The rationale behind this is to generate spotlights over summary with higher rewards. This results in the KTO dataset being twice the size of the original dataset, with an equal number of upvoted and downvoted examples. We then apply KTO to the instruction-tuned versions of base models (e.g., Llama-7b-chat-hf, Mistral-7B-Instruct-v0.2) using the KTO dataset, balancing desirable and undesirable weights. Under this setup, we evaluate the performance of the KTO on three LLMs (Mistral-7b, Llama2-7b, and Llama3-8b) by comparing its generations with ground truths on the test set. We also apply KTO on SFT but the results are not better. Among all experiments with different versions of base models, we report the best outcome only (which was achieved when KTO was done directly on the instruct models). These results are denoted in Tables 20, 22, 24, 26 in rows denoted by KTO(Mst-7b), KTO(L2-7b) and KTO(L3-Bb).
@@ Appendix E Related Work
1. Summarization:
Researchers focus on summarization task on different domains - news documents Du and Huo (2020); Goyal et al. (2022); Huang et al. (2014); Fikri et al. (2021), scientific articles Cohan and Goharian (2016, 2017); Cachola et al. (2020) and other different areas - social media Mukherjee et al. (2022); Chakraborty et al. (2019), wikipedia Ta et al. (2023); Frefel (2020) etc. TLDR Cachola et al. (2020), TLDR9+ Sotudeh et al. (2021), CiteSum Mao et al. (2022), Narayan et al. (2018) work on extreme summarization of main concepts of articles. People also focus on different summarization - query focused Vig et al. (2022); Baumel et al. (2016); Liu et al. (2024), layman summarization Goldsack et al. (2023); Salaün et al. (2022); Luo et al. (2022), goal oriented Goldstein et al. (2007); Ham et al. (2020); Jin et al. (2024) etc. CTRLSum He et al. (2020), Zhang et al. (2023) and  Ribeiro et al. (2023) focus on summaries with controllable readability. Researchers also work on summarizing documents with respect to key elements Ryu et al. (2024); Wang et al. (2023). But, the problem of spotlight generation differs from these summarization tasks.
2. Highlight Generation:
People focus on highlight detection task in different directions - story highlight Woodsend and Lapata (2010), research paper highlight Rehman et al. (2023a, b), micro-blog news highlight Wei and Gao (2018), Teaser generation Karn et al. (2019), highlight span extraction Arumae et al. (2019); Cho et al. (2020). But, these work mostly focus on extracting multiple spans as a highlight. Our works differ from the fact that we generate most important part of the article which will entice reader to read through the article.
3. Headline/Title Generation:
Researchers aim at developing different approaches - sequence prediction Colmenares et al. (2015) for headline generation, personalized headline generation Ao et al. (2021, 2023), self-attention based headline generation Gavrilov et al. (2019). But our task is different from traditional headline generation work which is too short.
4. Teaser Production:
Teasers serve as a powerful strategy for promoting content across various domains. Earlier works cover teaser generation from long documents Xu et al. (2024), news article Karn et al. (2019), ted-talks Vico and Niehues (2022) etc.
Our work differs from all of the above as Spotlight has different features than summary, highlight and headline.
@@ G.1 Task 1 - Identify Interest-Generation
Please do not use ChatGPT/GPT4 or any Large Language Models - all reponses should be generated by human properly. It is a strict instruction and will be checked manually - if any issue: the submission will be rejected and re-doing will be required.
Participant Prescreening criteria.
@@ G.2 Task 2 - Summary Generation
Carefully review the Source Document. No additional software download is required. Use a browser, preferably Google Chrome, and ensure a stable internet connection. Allocate time judiciously for crafting each of the summaries based on the provided 10 SD instances. After completion, you will be asked to provide feedback on the generation exercise, platform interaction, and details about your academic background, age, country of birth, and any medical background or experience with the summary generation task. Your summary length (word count) must be around 15% ($\pm$ 20 words) of the document. For example if the Source Document has 1000 words, the summary must have between 130 to 170 words. Please do not use ChatGPT/GPT4 or any Large Language Models - all reponses should be generated by human properly. It is a strict instruction and will be checked manually - if any issue: it will be rejected and re-doing will be required.
@@
G.3 Task 3 - Mini-Story Validation:
@@
Please do not use ChatGPT/GPT4 or any Large Language Models - all reponses should be generated by human properly. It is a strict instruction and will be checked manually - if any issue: the submission will be rejected and re-doing will be required.
Participant Prescreening criterias:-
@@ Appendix H Examples
One examples of human annotation interface is shown in Fig 5.
@@ I.1 Performance of Traditional Baseline Approaches
In addition to the results in Table 4, we have also explored different traditional approaches for spotlight generation and results for News, CSPubSum, Wikipedia and Research Presentation datasets are shown in Table 19, 21, 23 and 25. It is seen that even simple SFT-based approaches on LLMs perform much better than these traditional approaches across all metrics. This performance gap may be due to the advanced architectures and targeted optimizations of LLM models which are fine-tuned specifically for task performance and user preferences, whereas the older models do not benefit from the same level of refinement and cutting-edge techniques, leading to their comparatively lower performance.
@@ I.2 Performance of DPO vis-a-vis other LLM-based approaches
In addition to the results presented in Table 5, we demonstrate the effectiveness of the DPO and KTO approach for spotlight generation across various evaluation criteria for Llama3-8b (L3-8b), Llama2-7b (L2-7b) and Mistral-7b (Mst-7b) large language model (LLM) for each datasets in Table 20 (for News), 22 (for CSPubSum), 24 (for Wikipedia) and 26 (for Research Presentation). We report evaluations for Llama2-7b, Mistral-7b and Llama3-8b finetuning abbreviated as L2-7b-FT, Mst-7b-FT and L3-8b-FT, general instruction tuned vanilla model L2-7b-VA, Mst-7b-VA and L3-8b-VA, DPO DPO(L2-7b), DPO(Mst-7b) and DPO(L3-8b) and KTO KTO(L2-7b), KTO(Mst-7b) and KTO(L3-8b). It is seen that
@@
  • Across all the datasets, both Llama2 and Mistral models consistently show that DPO optimization leads to superior performance over KTO—evidenced by higher ROUGE, BERT-F1, UniEval relevance, and GPT-4 Critique scores. This aggregated trend reinforces the robustness of the DPO approach in spotlight generation across diverse domains.
  • The experimental results indicate that the DPO-based approach outperforms large LLMs (GPT-3.5, Gemini and GPT-4) in 1-shot evaluations for spotlight generation.
  • Also, to validate the GPT-4 critique scores against robust human judgment, we conducted an additional human annotation study similar to Mullick et al. (2024) to evaluate the alignment between human judgment and automated assessments for 20 documents for each dataset. In this study, annotators were presented with the model-generated spotlight and asked to categorize them into one of two options - ‘good’ or ‘bad’. Then, we calculate the Point-biserial correlation coefficient between GPT-4 critique scores and human judgment. We obtain an overall average of 0.751 point-biserial correlation coefficient between them, which validates GPT-4 critique scores are robust and reliable against human judgment studies.

### v3

@@ 7 Conclusion
In this paper, we introduced the concept of “Spotlight,” a novel, concise representation of source documents that offers a different perspective to conventional summaries and highlights. Our work identifies a set of unique characteristics that differentiate spotlights, emphasizing their ability to present key aspects of a document in a faithful yet compelling way, effectively capturing the reader’s attention as well as drawing their attention to the source document. To validate this concept, we prepared four carefully annotated datasets, which served as the basis for a comprehensive analysis. Through rigorous quantitative evaluations and human studies, we established nine key features that differentiate spotlights primarily in terms of coverage and presentation style. Building on these findings, we proposed a Direct Preference Optimization (DPO)-based fine-tuning algorithm to generate spotlights. Extensive evaluation using both traditional metrics and LLM-based assessments demonstrate the effectiveness of our approach. We believe that, similar to summarization, there is a significant potential for developing various types of spotlights — such as aspect-based, query-focused, and personalized versions — which will be the focus of our immediate future work.
@@
Our datasets are neither multilingual nor multimodal. So, we need to explore how state-of-the-art approaches can be utilized in multilingual and multimodal scenarios for spotlight generation - which we aim to do as a part of future work. We have also explored some edge cases and limitations of the model generated spotlights as in Appendix J.
@@ Appendix A Dataset
Figure 2: Examples of (Document, Spotlight) pairs from different datasets, An example from CSPubSum is given in Figure 1, The documents and the example spotlight of Research Presentation have been truncated in this visualization owing to their length. Across the three examples, the spotlights consistently transform lengthy, detail-heavy texts into short, curiosity-driven narratives. In the News case, procedural details and background information are removed, leaving a striking account that emphasizes the unusual and attention-grabbing aspect of the story—the body discovered after four days. The Wikipedia example shifts from an encyclopedic biographical entry to a narrative that foregrounds milestones and memorable achievements, making the article more engaging and easier to recall. Finally, the Research Presentation example reframes dense, technical exposition into an accessible narrative by highlighting the central research question and its intrigue rather than methodological detail. Together, these examples illustrate how spotlights distill complex information into focused, compelling narratives that spark interest rather than merely summarize content.
@@
Figure 3: News Headline and the Short Description taken as spotlight for the News dataset. This figure illustrates the construction process of the News spotlight dataset. Source articles are paired with human-curated title and short description.
Figure 4: Author written highlights taken as spotlight for the CSPubSum dataset. From the Research Papers the author-written bullet points serve as a compelling entryway into the research paper, offering a concise preface that draws attention to the central questions and findings while setting an engaging narrative tone.
Figure 5: Conference Video Transcription taken as spotlight for the Research Presentation dataset. From slide decks and talk transcripts, concise spotlights are crafted to reflect the most engaging aspects of the research story. Unlike abstracts or slide summaries, these spotlights focus on intrigue, contributions, and memorable phrasing, ensuring that the dataset captures narrative appeal specific to academic communication.
@@ B.1 Readability
Table 11: Feature level analysis of spotlight (Spot) vs. human annotated summary (Summ) for 60 articles of each datasets, For Entropy, Compactness, Flesch Kincaid, Gunning score, Dale-Chall and McLaughlin’s score lower values signifies better result, whereas for Skew50, Flesch Reading score, FORCAST and Interest-Generation higher values signifies a better result as discussed in Section 3, More Exploratory Results of Table 3.
@@ B.2 Informativeness and Degree of Extraction
We analyze the number of sentences in the document that achieve the R1, R2, and RL score above a specified threshold th with at least one sentence in the summary or spotlight. This approach assumes that if a sentence meets the threshold, its information has been incorporated into the summary or spotlight - i.e. informativeness. However, sentences incorporate information in an abstractive fashion so the threshold th needs to be set at a relatively low value. At a value of $th=0.3$ our findings of different Rouge (1, 2 and L) scores at the Column ‘Informativeness’ in Table 12 and 14 respectively for human-generated and GPT-4 generated summaries, indicates that a larger number of sentences from the source document are incorporated into summaries compared to spotlights. It signifies that the summary is more informative than the spotlight.
However, when the threshold value is increased (at a value of $th=0.8$) — indicating a higher level of extractiveness — the trend reverses, with fewer sentences meeting the threshold in summaries compared to spotlights as shown in ‘Degree of Extraction’ column using different Rouge (1, 2 and L) scores in Table 12 and 14 respectively for human-generated and GPT-4 generated summaries. The increased extractiveness of the spotlight can also be attributed to its enhanced readability. This trend is consistent across different ROUGE metrics (R1, R2, RL), as shown in the mentioned Tables.
@@
B.3 Interest Generation aspects of a spotlight
@@
Figure 6: Interest-Generation aspect of a spotlight on Life and Career of Anja Nissen from Wikipedia dataset, where annotators selectively identify the most engaging and narratively significant parts of a text. The bold segments represents the achievements, competitions and performances of Anja Nissen that capture attention and convey intrigue, rather than covering every detail. This mirrors the idea of spotlight generation which focus on a few compelling elements that can spark curiosity and draw readers into the broader narrative.
@@
Dataset Both of them Spotlight only Summary only None of them News 6 34 6 4 CSPubSum 11 32 4 3 Wikipedia 12 31 1 6 Research 11 30 1 8 Presentation
@@ Appendix C Spotlight Generation Process
KTO Ethayarajh et al. (2024): We initially use the summaries for documents in the dataset to create a binary preference Kahneman-Tversky Optimization (KTO) dataset. Each document from the original dataset contributes one upvoted and one downvoted example. The document paired with the ground truth spotlight forms the upvoted example (indicating a preferred completion), while the document paired with a summary forms the downvoted example (denoting undesirable). The rationale behind this is to generate spotlights over summary with higher rewards. This results in the KTO dataset being twice the size of the original dataset, with an equal number of upvoted and downvoted examples. We then apply KTO to the instruction-tuned versions of base models (e.g., Llama-7b-chat-hf, Mistral-7B-Instruct-v0.2) using the KTO dataset, balancing desirable and undesirable weights. Under this setup, we evaluate the performance of the KTO on three LLMs (Mistral-7b, Llama2-7b, and Llama3-8b) by comparing its generations with ground truths on the test set. We also apply KTO on SFT but the results are not better. Among all experiments with different versions of base models, we report the best outcome only (which was achieved when KTO was done directly on the instruct models). These results are denoted in Tables 21, 23, 25, 27 in rows denoted by KTO (Mst-7b), KTO (L2-7b) and KTO (L3-Bb).
@@ Appendix E Related Work
1. Summarization: Researchers focus on summarization task on different domains - news documents Du and Huo (2020); Goyal et al. (2022); Huang et al. (2014); Fikri et al. (2021), scientific articles Cohan and Goharian (2016, 2017); Cachola et al. (2020) and other different areas - social media Mukherjee et al. (2022); Chakraborty et al. (2019); Mullick et al. (2024d), wikipedia Ta et al. (2023); Frefel (2020) , dialog Mullick et al. (2024a) etc. TLDR Cachola et al. (2020), TLDR9+ Sotudeh et al. (2021), CiteSum Mao et al. (2022), Narayan et al. (2018) work on extreme summarization of main concepts of articles. People also focus on different summarization - query focused Vig et al. (2022); Baumel et al. (2016); Liu et al. (2024), layman summarization Goldsack et al. (2023); Salaün et al. (2022); Luo et al. (2022), goal oriented Goldstein et al. (2007); Ham et al. (2020); Jin et al. (2024), persona-based Mullick et al. (2024c) etc. CTRLSum He et al. (2020), Zhang et al. (2023) and  Ribeiro et al. (2023) focus on summaries with controllable readability. Researchers also work on summarizing documents with respect to key elements Ryu et al. (2024); Wang et al. (2023). But, the problem of spotlight generation differs from these summarization tasks.
2. Highlight Generation: People focus on highlight detection task in different directions - story highlight Woodsend and Lapata (2010), research paper highlight Rehman et al. (2023a, b), micro-blog news highlight Wei and Gao (2018), Teaser generation Karn et al. (2019), highlight span extraction Arumae et al. (2019); Cho et al. (2020). But, these work mostly focus on extracting multiple spans as a highlight. Our works differ from the fact that we generate most important part of the article which will entice reader to read through the article.
3. Headline/Title Generation: Researchers aim at developing different approaches - sequence prediction Colmenares et al. (2015) for headline generation, personalized headline generation Ao et al. (2021, 2023), self-attention based headline generation Gavrilov et al. (2019). But our task is different from traditional headline generation work which is too short.
4. Teaser Production: Teasers serve as a powerful strategy for promoting content across various domains. Earlier works cover teaser generation from long documents Xu et al. (2024), news article Karn et al. (2019), ted-talks Vico and Niehues (2022) etc.
Our work differs from all of the above as Spotlight has different features than summary, highlight, headline and teaser, and serves different perspectives. This concept of spotlight is also different from the notion of entity-extraction Mullick et al. (2024e); Guha et al. (2021); Mullick et al. (2022c, 2021), opinion context detection Mullick et al. (2017b, 2016, 2018a, 2018b, 2019, a). intent phrase identification task in specialized contexts Mullick et al. (2022d, b); Mullick (2023b, a); Mullick et al. (2022a, 2024b, 2023, 2024f).
@@ G.1 Task 1 - Identify Interest-Generation
Please do not use ChatGPT/GPT4 or any Large Language Models - all responses must be produced by human annotators. It is a strict instruction and will be checked manually - if any issue: the submission will be rejected and re-doing will be required.
Participant Prescreening criteria:
@@ G.2 Task 2 - Summary Generation
Carefully review the Source Document. No additional software download is required. Use a browser, preferably Google Chrome, and ensure a stable internet connection. Allocate time judiciously for crafting each of the summaries based on the provided 10 SD instances. After completion, you will be asked to provide feedback on the generation exercise, platform interaction, and details about your academic background, age, country of birth, and any medical background or experience with the summary generation task. Your summary length (word count) must be around 15% ($\pm$ 20 words) of the document. For example if the Source Document has 1000 words, the summary must have between 130 to 170 words. Please do not use ChatGPT/GPT4 or any Large Language Models - all reponse should be generated by human properly. It is a strict instruction and will be checked manually - if any issue: it will be rejected and re-doing will be required.
@@
G.3 Task 3 - Mini-Story Validation
@@
Please do not use ChatGPT/GPT4 or any Large Language Models - all reponse should be generated by human properly. It is a strict instruction and will be checked manually - if any issue: the submission will be rejected and re-doing will be required.
Participant Prescreening criteria:-
@@ Appendix H Examples
One example of the human annotation interface is shown in Fig 7.
@@ Appendix I Experimental Results
We analyzed the performance of traditional and LLM-based approaches, including the preference optimization techniques using various referenceless metrics and LLM-critiquing. We also assess different variations of DPO with held-out data.
@@ I.1 Performance of Traditional Baseline Approaches
In addition to the results in Table 5, we have also explored different traditional approaches for spotlight generation and results for News, CSPubSum, Wikipedia and Research Presentation datasets are shown in Table 20, 22, 24 and 26. It is seen that even simple SFT-based approaches on LLMs perform much better than these traditional approaches across all metrics. This performance gap may be due to the advanced architectures and targeted optimizations of LLM models which are fine-tuned specifically for task performance and user preferences, whereas the older models do not benefit from the same level of refinement and cutting-edge techniques, leading to their comparatively lower performance.
@@ I.2 Performance of DPO vis-a-vis other LLM-based approaches
In addition to the results presented in Table 6, we demonstrate the effectiveness of the DPO and KTO approach for spotlight generation across various evaluation criteria for Llama3-8b (L3-8b), Llama2-7b (L2-7b) and Mistral-7b (Mst-7b) large language model (LLM) for each datasets in Table 21 (for News), 23 (for CSPubSum), 25 (for Wikipedia) and 27 (for Research Presentation). We report evaluations for Llama2-7b, Mistral-7b and Llama3-8b finetuning abbreviated as L2-7b-FT, Mst-7b-FT and L3-8b-FT, general instruction tuned vanilla model L2-7b-VA, Mst-7b-VA and L3-8b-VA, DPO DPO(L2-7b), DPO(Mst-7b) and DPO(L3-8b) and KTO KTO(L2-7b), KTO(Mst-7b) and KTO(L3-8b). It is seen that
@@
  • Across all the datasets, both Llama2 and Mistral models consistently show that DPO optimization leads to superior performance over KTO—evidenced by higher ROUGE, BERT‑F1, UniEval relevance, and GPT‑4 Critique scores. This aggregated trend reinforces the robustness of the DPO approach in spotlight generation across diverse domains.
  • The experimental results indicate that the DPO-based approach outperforms large LLMs (GPT-3.5, Gemini and GPT-4) in 1‑shot evaluations for spotlight generation.
  • Also, to validate the GPT-4 critique scores against robust human judgment, we conducted an additional human annotation study similar to Mullick et al. (2024c) to evaluate the alignment between human judgment and automated assessments for 20 documents for each dataset. In this study, annotators were presented with the model-generated spotlight and asked to categorize them into one of two options - ‘good’ or ‘bad’. Then, we calculate the Point-biserial correlation coefficient between GPT-4 critique scores and human judgment. We obtain an overall average of 0.751 point-biserial correlation coefficient between them, which validates GPT-4 critique scores are robust and reliable against human judgment studies.